OpenStudy (anonymous):

Evaluate the expression by using the given value of the variables. y ÷ 2 • 4 – 10 ÷ x + 6 where x = 5 and y = 16 A. 3 B. 21 C. –3 D. 36

OpenStudy (anonymous):

i got 21 is it right

OpenStudy (anonymous):

wat about this question Solve for y. 2y – 16 = –4

OpenStudy (kropot72):

For the first question the operations must be done in the following order: 1. Division 2. Multiplication 3. Addition 4. Subtraction 21 is not the correct answer. Can you try again using the correct order of operations?

OpenStudy (anonymous):

ill try thx

OpenStudy (kropot72):

Come back if you need more help.

OpenStudy (anonymous):

hmm i got 20

OpenStudy (kropot72):

\[16\div 2\times 4 -10\div 5+6\] Carrying out division gives: \[8\times 4 -2 +6\] Carrying out multiplication gives: 32 - 2 + 6 = you can find the answer.

OpenStudy (anonymous):

thx

OpenStudy (kropot72):

You're welcome :)

OpenStudy (anonymous):

:)